Status Report

NOAA SATOPS Morning Report: Friday, June 22, 2012

By SpaceRef Editor
June 22, 2012
Filed under ,

SATOPS Morning Report: Friday, June 22, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 174

GOES
GOES-13 DOY173 1750Z – Engineering commanded the Sounder back to a single motor winding (#2) operation. Unexpected scan line length and data dropouts immediately appeared as observed during the initial anomaly. At 1758Z the Sounder was commanded back to utilizing a dual motor winding configuration. The instrument continues to be assessed and the plan is to remain on both windings until further notice. The 1746Z Sounding was degraded as a result of the activity.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
Nothing reported.

Questions or comments to [email protected]

SpaceRef staff editor.